A man of mass m1= 70.0 kg is skating at v1= 8.00 m/s behind his wife of mass m2= 50.0 kg, who is skating at v2=4.00 m/s. Instead of passing her, he inadvertently collides with her. He grabs her around the waist, and they maintain their balance. (a) Sketch the problem with before - and - after diagrams, representing the skaters as blocks. (b) Is the collision best described as elastic, inelastic, or perfectly inelastic? Why? (c) Write the general equation for conservation of momentum in terms of m1, v1, m2, v2, and final velocity vf . (d) Solve the momentum equation for vf

1
Expert's answer
2021-05-26T17:22:24-0400

(a) Let's draw before- and after diagrams:



(b) Collision is perfectly inelastic, because it occurs when the maximum amount of kinetic energy of a system is lost. As a result, there is a zero coefficient of restitution and colliding bodies stick together.

(c) From the lwa of conservation of momentum, we have:


"m_1v_1+m_2v_2=(m_1+m_2)v_f."

(d)

"v_f=\dfrac{m_1v_1+m_2v_2}{m_1+m_2},""v_f=\dfrac{70\ kg\cdot8\ \dfrac{m}{s}+50\ kg\cdot4\ \dfrac{m}{s}}{70\ kg+50\ kg}=6.3\ \dfrac{m}{s}."